OpenStudy (anonymous):

9x – 6y = 30 wouldnt it be y=30-9x/6? I put that in and it was wrong and i reduced it to y=5-9x/6 but it says its still wrong. Need some help please

OpenStudy (anonymous):

you forget the sign on the Y it's -6y

OpenStudy (anonymous):

9x - 6y = 30 -6y = -9x + 30 y = 3/2 - 5

OpenStudy (anonymous):

y=-(30-9x)/6

OpenStudy (anonymous):

oh ok thank you